OpenStudy (anonymous):

prove (1-sin x)/sec x = cos^3 x/1+sin x any trig people in the houseeeeee

OpenStudy (anonymous):

i started by multiplying the left side by the conjugate. but i know you can factor the other side too and start that way

OpenStudy (kmeis002):

With trig id problems, its helpful to start working on a side that has a lot of algebra that can be done. for example, I'll get your started \[ \frac{1-\sin x}{\sec x} = \frac{\cos^3 x}{1+\sin x} \] It seems that the right hand side has a lot of manipulation potential. we can: \[ \frac{1-\sin x}{\sec x} = \frac{(\cos^2 x)(\cos x)}{1+\sin x} \] Recall \[\cos^2x = 1-\sin^2 x \] Then \[ \frac{1-\sin x}{\sec x} = \frac{(1-\sin^2x)(\cos x)}{1+\sin x} \] Maybe you can multipled the right hand side by the conjugate to keep it going.

OpenStudy (kmeis002):

In fact, if you do, you may almost be done.

OpenStudy (anonymous):

no i dont know what im doing

OpenStudy (anonymous):

\[\huge \frac{ (1-\sin x)}{\sec x}\] multiplying Nr. and Dr. by (1+ sin x) we find \[\huge \frac{ (1-\sin x)}{\sec x} = \frac{ (1-\sin x)}{\sec x} \times \frac{ (1+\sin x)}{(1+\sin x)}\] \[\huge = \frac{ (1^2-\sin^2 x)}{\frac{1}{\cos x}(1+\sin x)}\] \[\huge = \frac{ \cos^2 x \times \cos x}{(1+\sin x)} = \frac{\cos^3x}{ 1 +sinx}\]

OpenStudy (anonymous):

I GET ITTTTTTTTTTTTTTTTTTTTTTT

OpenStudy (anonymous):

if you dont know those identities youre done. you fail.

OpenStudy (anonymous):

what happened with 1/cos x(1+sinx) for the demoninators

OpenStudy (kmeis002):

\[\sec x = \frac{1}{cosx} \] then simply by multiplying by the reciprocal

OpenStudy (anonymous):

how did that become 1+sinx

OpenStudy (johnweldon1993):

The (1 + sin(x) stayed but the 1/cos(x) was flipped and then multiplied to the fraction (that's where the 2nd cos came from in the numerator up there)

OpenStudy (kmeis002):

\[\frac{\cos^2 x}{\frac{1}{\cos x}(1+\sin x)} = \frac{\cos^2 x}{(1+\sin x)} \frac{\cos x}{1}\]

OpenStudy (johnweldon1993):

\[\huge \frac{ (1^2-\sin^2 x)}{\frac{1}{\cos x}(1+\sin x)} = \huge \frac{ (1^2-\sin^2 x)}{(1+\sin x)} \times cos(x)\]

OpenStudy (anonymous):

okay all of you have helped me so much i wish i could medal all of you